# A Survey on 3D Gaussian Splatting

### v8

@@ 4.7. New Rendering Algorithm for 3D Gaussians
While the rasterization-based pipeline of 3D GS offers impressive real-time performance, it still suffers from the inherent limitations, including inefficient handling of highly-distorted cameras (crucial for robotics), secondary rays (for optical effects like reflections and shadows), and stochastic ray sampling (needed in various existing pipelines). In addition, the assumptions that Gaussians do not overlap and can be sorted accurately using only centers are often violated in practice, leading to temporal artifacts when camera movement changes sorting order.
@@ 5. Application Areas and Tasks
Building on the rapid advancements in 3D GS, a wide range of innovative applications has emerged across multiple domains (c.f. Fig. 6) such as robotics (Sec. 5.1), dynamic scene reconstruction and representation (Sec. 5.2), generation and editing (Sec. 5.3), avatar (Sec. 5.4), medical systems (Sec. 5.5), large-scale scene reconstruction (Sec. 5.6), physics (Sec. 5.7), and even other scientific disciplines (Zhang et al., 2024j; Wu et al., 2024f; Zhang et al., 2024e; Li et al., 2024d). Here, we highlight key examples that underscore the transformative impact and potential of 3D GS and offer a more comprehensive collection in Github. We also provide a collection of representative datasets (c.f. Table 5.7) according to our taxonomy.
5.15.25.35.45.55.65.7(Yan et al., 2024c), (Keetha et al., 2024), (Matsuki et al., 2024), (Yugay et al., 2023), (Huang et al., 2024c), (Li et al., 2024e), (Ji et al., 2024a)(Lu et al., 2024c), (Abou-Chakra et al., 2024), (Shorinwa et al., 2024), (Ji et al., 2024b), (Zheng et al., 2024a), (Zhu et al., 2024b), (Peng et al., 2024a)(Luiten et al., 2024), (Yang et al., 2024b), (Wu et al., 2024c), (Yang et al., 2024d), (Yu et al., 2024b), (Shao et al., 2024b)(Huang et al., 2024d), (Das et al., 2024), (Lin et al., 2024a), (Li et al., 2024a), (Sun et al., 2024a), (Lu et al., 2024a)(Tang et al., 2024b), (Yi et al., 2024), (Tang et al., 2024a), (Zhou et al., 2024b), (Li et al., 2024b), (Mu et al., 2024)(Jiang et al., 2024a), (Wang et al., 2024a), (Chen et al., 2024a), (Fang et al., 2024), (Zhou et al., 2024a), (Qiu et al., 2024)(Li et al., 2024k), (Hu and Liu, 2024), (Lei et al., 2024a), (Yuan et al., 2024)(Zhou et al., 2024d), (Qian et al., 2024a), (Dhamo et al., 2023), (Li et al., 2024j)(Huang et al., 2024b), (Liu et al., 2024b), (Zhu et al., 2024c)(Zhao et al., 2024d), (Wang et al., 2024c), (Bonilla et al., 2024)(Wu et al., 2024g), (Wu et al., 2024b), (Xiong et al., 2024), (Zhao et al., 2024c)(Cheng et al., 2024), (Kerbl et al., 2024), (Liu et al., 2024a), (Lin et al., 2024b), (Ren et al., 2024)(Xie et al., 2024), (Liu et al., 2024e), (Borycki et al., 2024), (Qiu et al., 2024), (Huang et al., 2025)(Zhang et al., 2024h), (Jiang et al., 2024c), (Feng et al., 2024a), (Zhong et al., 2024), (Shao et al., 2024a)
Figure 6. Typical applications benefited from GS (Sec. 5). Some images are borrowed from (Lu et al., 2024a; Tang et al., 2024a; Zhu et al., 2024c; Xiong et al., 2024; Lei et al., 2024a; Xie et al., 2024) and redrawn.
@@ 5.1. Robotics
The integration of GS into robotic systems has yielded significant advancements across three core domains. In SLAM, GS-based methods (Huang et al., 2024c; Matsuki et al., 2024; Yugay et al., 2023; Yan et al., 2024c; Keetha et al., 2024; Hong et al., 2024; Ji et al., 2024a; Sun et al., 2024b; Tosi et al., 2024; Li et al., 2024e; Zhu et al., 2024b; Deng et al., 2024; Hu et al., 2024a; Lang et al., 2024; Peng et al., 2024a) excel in real-time dense mapping but face inherent trade-offs. Visual SLAM frameworks, particularly RGB-D variants (Keetha et al., 2024; Yugay et al., 2023; Sun et al., 2024b), leverage depth supervision for geometric fidelity but falter in low-texture or motion-degraded environments. RGB-only approaches (Matsuki et al., 2024; Huang et al., 2024c; Sandström et al., 2024) circumvent depth sensors but grapple with scale ambiguity and drift. Multi-sensor fusion strategies, such as LiDAR integration (Hong et al., 2024; Lang et al., 2024; Wu et al., 2024b), enhance robustness in unstructured settings at the cost of calibration complexity. Semantic SLAM (Zhu et al., 2024b; Ji et al., 2024a; Li et al., 2024e) extends scene understanding through object-level semantics but struggles with scalability due to lighting sensitivity in color-based methods or computational overhead in feature-based methods. 3D GS based manipulation (Lu et al., 2024c; Abou-Chakra et al., 2024; Shorinwa et al., 2024; Ji et al., 2024b; Zheng et al., 2024a) bypasses the need for auxiliary pose estimation in NeRF-based methods, enabling rapid single-stage tasks like grasping in static environments via geometric and semantic attributes encoded in Gaussian properties. Multi-stage manipulation (Lu et al., 2024c; Shorinwa et al., 2024), where environmental dynamics demand real-time map updates, requires explicit modeling of dynamic adjustments (e.g., object motions and interactions), material compliance, etc.
@@ 5.2. Dynamic Scene Reconstruction
Dynamic scene reconstruction refers to the process of capturing and representing the three-dimensional structure and appearance of a scene that changes over time (Pumarola et al., 2021; Park et al., 2021b, a; Guo et al., 2023). This involves creating a digital model that accurately reflects the geometry, motion, and visual aspects of the objects in the scene as they evolve. Dynamic scene reconstruction is crucial in various applications, e.g., VR/AR, 3D animation, and autonomous driving (Zhou et al., 2024c; Yan et al., 2024a; Zhou et al., 2024e).
@@
The key to adapt 3D GS to dynamic scenes is the modeling of temporal dimension which allows for the representation of scenes that change over time. 3D GS based methods (Yang et al., 2024b, d; Lin et al., 2024a; Kratimenos et al., 2023; Das et al., 2024; Shaw et al., 2023; Wu et al., 2024c; Luiten et al., 2024; Huang et al., 2024d; Shao et al., 2024b; Yu et al., 2024b; Liang et al., 2023; Katsumata et al., 2023; Li et al., 2024a; Guo et al., 2024; Bae et al., 2024; Lei et al., 2024b; Wang et al., 2024d; Duan et al., 2024) for dynamic scene reconstruction can generally be divided into two main categories as discussed in Sec. 4.5 and Sec. 4.6. The first category utilizes additional fields like spatial MLPs or grids to model deformation (Sec. 4.6). For example, Yang et al. (Yang et al., 2024b) first proposed deformable 3D Gaussians tailored for dynamic scenes. These 3D Gaussians are learned in a canonical space and can be used to model spatial-temporal deformation with an implicit deformation field (implemented as an MLP). GaGS (Lu et al., 2024a) devised the voxelization of a set of Gaussian distributions, followed by the use of sparse convolutions to extract geometry-aware features, which are then utilized for deformation learning. On the other hand, the second category is based on the idea that scene changes can be encoded into the 3D Gaussian representation with a specially designed rendering process (Sec. 4.5). For instance, Luiten et al. (Luiten et al., 2024) introduced dynamic 3D Gaussians to model dynamic scenes by keeping the properties of 3D Gaussians unchanged over time while allowing their positions and orientations to change. Yang et al. (Yang et al., 2024d) designed a 4D Gaussian representation, where additional properties are used to represent 4D rotations and spherindrical harmonics, to approximate the spatial-temporal volume of scenes.
@@
While 3D GS advances dynamic scene reconstruction by modeling per-Gaussian deformations, its reliance on fine-grained primitives limits scalability and robustness. Current methods struggle to balance computational efficiency and precision: small-scale reconstructions unify dynamic and static elements but become intractable in large environments, often requiring manual priors to segment regions —– a barrier in unstructured settings. Furthermore, the absence of object-level motion reasoning leads to artifacts and poor generalization over long sequences. Future work might prioritize object-centric frameworks that group Gaussians into persistent entities so as to model inherent motion disentanglement (dynamic vs. static).
@@ 5.3. Generation and Editing
Content generation and editing represent two fundamental and inherently interconnected capabilities in modern AI systems. While generation enables the synthesis of novel digital content from scratch or conditional inputs (Goodfellow et al., 2020; Ho et al., 2020; Rombach et al., 2022), editing provides the crucial ability to refine, adapt, and manipulate existing content with precise control (Zhang et al., 2023a). Together, these capabilities revolutionize creative workflows by combining initial content creation with iterative refinement, enabling applications from professional content production to interactive consumer tools.
Recent advances in generation (Chen et al., 2024b; Tang et al., 2024b; Yi et al., 2024; Liang et al., 2024a; Liu et al., 2024f; Yang et al., 2023a; Zou et al., 2024; Ling et al., 2024; Ren et al., 2023; Yin et al., 2023; Zhang et al., 2023b; Pan et al., 2024; Xu et al., 2024c; Tang et al., 2024a; Yang et al., 2024c; Barthel et al., 2024; Jiang and Wang, 2024; Zhou et al., 2024b; Zhuo et al., 2024; Wu et al., 2024d; He et al., 2024; Yang and Wang, 2024; Kim et al., 2024; Feng et al., 2024b; Li et al., 2024g; Melas-Kyriazi et al., 2024; Li et al., 2024b; Zhang et al., 2024b; Mu et al., 2024; Lee et al., 2024a) have led to the emergence of three main approaches. Optimization based methods (Chen et al., 2024b; Tang et al., 2024b; Yi et al., 2024) distill diffusion priors (gradients) to guide 3D model updates with the score functions. While these methods demonstrate impressive fidelity, they face significant computational overhead due to the necessity of comparing multiple viewpoints during the optimization process. Reconstruction based methods (Tang et al., 2024a; Lee et al., 2024a; Melas-Kyriazi et al., 2024) reframe the generation problem as a multi-view reconstruction task utilizing pre-trained multi-view diffusion models. Although this approach offers an intuitive and straightforward solution, it grapples with fundamental limitations in maintaining view consistency. The lack of strict geometric constraints across different viewpoints often results in inconsistent surface geometry and degraded texture quality, particularly in regions with complex visual features. Direct 3D generation methods train diffusion models on 3D representations (Mu et al., 2024; He et al., 2024; Zhang et al., 2024b). While the learned 3D diffusion models facilitate multi-view consistency, the demanding computational costs impede the expansion of training scales necessary for improved generative diversity.
Current editing works (Chen et al., 2024a; Fang et al., 2024; Huang and Yu, 2023; Ye et al., 2024; Cen et al., 2023; Zhou et al., 2024a; Lan et al., 2023; Huang et al., 2024d; Shao et al., 2024b; Yu et al., 2024b; Zhuang et al., 2024; Dou et al., 2024; Hu et al., 2024b; Palandra et al., 2024; Gu et al., 2024; Lyu et al., 2024; Liu et al., 2024c; Qiu et al., 2024; Zhang et al., 2024a; Wang et al., 2024a; Zhang et al., 2024g; Wu et al., 2024a) fall into two primary classes. The first class leverages 2D image-editing models (e.g., diffusion-based editors) to iteratively refine 3D Gaussians. Early efforts (Chen et al., 2024a; Fang et al., 2024; Palandra et al., 2024) adopt optimization- or reconstruction-based strategies akin to methods in generation, but introduce task-specific control signals. However, naively applying 2D edits independently across views often introduces multi-view inconsistencies. Subsequent works (Wang et al., 2024e, a; Zhang et al., 2024g; Wu et al., 2024a) mitigate this through iterative refinement or cross-view attention, albeit at increased computational costs for alignment. A notable challenge is unintended object deformations, attributed to the weak 3D geometric priors in 2D editing models and the difficulty of reconciling 2D edits with underlying 3D structures. The second class exploits the explicit nature of 3D GS to enable direct manipulation based on embedded properties such as semantics (Hu et al., 2024b; Cen et al., 2023; Ye et al., 2024; Qiu et al., 2024) and key points (Huang et al., 2024d). However, this class remains underexplored due to essential challenges: the lack of inherent ordering complicates the design of efficient indexing schemes, while editing attributes (e.g., texture and geometry) requires careful regularization and alignment to preserve plausibility.

### v9

@@ 4.7. New Rendering Algorithm for 3D Gaussians
While the rasterization-based pipeline of 3D GS offers impressive real-time performance, it still suffers from the inherent limitations, including inefficient handling of highly-distorted cameras (for robotics), secondary rays (for optical effects like reflections and shadows), and stochastic ray sampling (needed in various existing pipelines). In addition, the assumptions that Gaussians do not overlap and can be sorted accurately using only centers are often violated in practice, leading to temporal artifacts when camera movement changes sorting order. This issue is closely related to the long-standing problem of transparency and visibility ordering in classical rasterization, where order-independent transparency (OIT) methods aim to reduce sorting artifacts. Weighted-blended OIT (McGuire and Bavoil, 2013) provides a widely-used solution that approximates transparency without explicit per-pixel sorting.
@@
4.8. Kernel Engineering
Gaussian kernel is intrinsically smooth with infinite spatial support. Consequently, representing sharp boundaries, thin surfaces, and piecewise-planar structures can be parameter-inefficient: capturing high-frequency details requires aggressive densification and careful regularization. Furthermore, overlapping semi-transparent splats can produce cloudy artifacts near discontinuities due to the kernel’s low-pass nature and cross-boundary leakage.
To address these limitations, a growing body of work revisits the geometric primitive and the splatting operator. These efforts can be broadly categorized as follows: i) Surface-aligned primitives (Huang et al., 2024a; Dai et al., 2024) constrain splatted elements to align with local tangent planes. By utilizing planar or near-planar parameterizations alongside geometric regularizers, this direction explicitly forces primitives to match the underlying surface geometry, thereby improving multi-view consistency and the quality of extracted meshes. ii) Alternative distributions and shape modifications (Li et al., 2025; Huang et al., 2025b; Zhu et al., 2025; Liu et al., 2025b, a) alter the mathematical formulation and spatial footprint of the kernel to handle sharp transitions. By replacing the standard Gaussian with bounded-support Beta distributions, asymmetric cuts, deformable radial bases, or heavy-tailed mixtures with subtractive components (e.g., scooping), these methods provide finer control over local attenuation. This prevents cross-boundary density leakage and allows for a more compact representation of complex frequency profiles and hard edges. iii) Non-Gaussian geometric primitives and operator redesigns (Held et al., 2025; Franke et al., 2024; Ye et al., 2025; Condor et al., 2025) abandon the standard ellipsoid in favor of alternative shapes, or fundamentally modify the rendering operator (Sec. 4.7). This includes utilizing 3D smooth convexes, trilinear point splatting into image pyramids, or bi-scale pipelines that combine standard graphics rasterization with sorting-free splatting. These redesigns natively accommodate distinct geometric features and bypass the limitations of pure Gaussian rasterization.
Despite rapid progress, kernel engineering introduces new practical trade-offs. Alternative parameterizations consistently require custom rendering implementations and tailored primitive management strategies (e.g., modified splitting, merging, and pruning heuristics). Furthermore, standardized evaluation remains challenging, as comparisons are frequently confounded by differences in primitive budgets. As early hybrid pipelines begin to emerge, a promising direction lies in the principled compression and adaptive allocation of heterogeneous primitive types – deploying computationally complex kernels only where specific geometric intricacies demand them.
@@ 5. Application Areas and Tasks
Building on the rapid advancements in 3D GS, a wide range of innovative applications has emerged across multiple domains (c.f. Fig. 3) such as robotics (Sec. 5.1), dynamic scene reconstruction and representation (Sec. 5.2), generation and editing (Sec. 5.3), avatar (Sec. 5.4), medical systems (Sec. 5.5), large-scale scene reconstruction (Sec. 5.6), physics (Sec. 5.7), and even other scientific disciplines (Zhang et al., 2024h; Wu et al., 2024e; Zhang et al., 2024i; Li et al., 2024a). Here, we highlight key examples that underscore the transformative impact and potential of 3D GS and offer a more comprehensive collection in Github. We also provide a collection of representative datasets (c.f. Table Supplementary Material) according to our taxonomy.
5.15.25.35.45.55.65.7(Yan et al., 2024a; Keetha et al., 2024; Matsuki et al., 2024)(Lu et al., 2024a; Shorinwa et al., 2024)(Yang et al., 2024d)(Luiten et al., 2024; Yang et al., 2024b)(Tang et al., 2024b)(Tang et al., 2024a)(Mu et al., 2024; He et al., 2024; Pan et al., 2024a)(Chen et al., 2024c; Fang et al., 2024; Palandra et al., 2024)(Huang et al., 2024d; Waczyńska et al., 2024; Gao et al., 2024)(Li et al., 2024j; Lei et al., 2024a)(Qian et al., 2024a; Zhou et al., 2024e; Li et al., 2024d)(Huang et al., 2024e; Zhao et al., 2024a)(Wang et al., 2024b)(Liu et al., 2024d; Lin et al., 2024a)(Kerbl et al., 2024)(Xie et al., 2024; Feng et al., 2024b; Borycki et al., 2024)(Huang et al., 2025a; Zhang et al., 2024j; Zhong et al., 2024)
Figure 3. Typical applications benefited from GS (Sec. 5). Some images are borrowed from (Lu et al., 2024c; Tang et al., 2024a; Zhu et al., 2024a; Xiong et al., 2024; Lei et al., 2024a; Xie et al., 2024) and redrawn.
@@ 5.1. Robotics
The integration of GS into robotic systems has yielded advancements across three core domains. In SLAM, GS-based methods (Huang et al., 2024b; Matsuki et al., 2024; Yugay et al., 2023; Yan et al., 2024a; Keetha et al., 2024; Hong et al., 2024; Ji et al., 2024c; Sun et al., 2024; Tosi et al., 2024; Li et al., 2024g; Zhu et al., 2024b; Deng et al., 2024; Gao et al., 2025, 2026; Feng et al., 2025; Hu et al., 2024a; Lang et al., 2024; Peng et al., 2024b) excel in real-time dense mapping but face inherent trade-offs. Visual SLAM frameworks, particularly RGB-D variants (Keetha et al., 2024; Yugay et al., 2023; Sun et al., 2024), leverage depth supervision for geometric fidelity but falter in low-texture or motion-degraded environments. RGB-only approaches (Matsuki et al., 2024; Huang et al., 2024b; Sandström et al., 2024) circumvent depth sensors but grapple with scale ambiguity and drift. Multi-sensor fusion strategies, such as LiDAR integration (Hong et al., 2024; Lang et al., 2024; Wu et al., 2024a), enhance robustness in unstructured settings at the cost of calibration complexity. Semantic SLAM (Zhu et al., 2024b; Ji et al., 2024c; Li et al., 2024g; Yan et al., 2026) extends scene understanding through object-level semantics but struggles with scalability due to lighting sensitivity in color-based methods or computational overhead in feature-based methods. GS-based SLAM can be viewed as the online counterpart of pose-free reconstruction in Sec. 4.4: both address unknown poses by jointly optimizing camera trajectories and the Gaussian map, but SLAM operates sequentially with real-time constraints and additional mechanisms to control drift (e.g., keyframe selection and loop closure). 3D GS based manipulation (Lu et al., 2024a; Abou-Chakra et al., 2024; Shorinwa et al., 2024; Ji et al., 2024a; Zheng et al., 2024b) bypasses the need for auxiliary pose estimation in NeRF-based methods, enabling rapid single-stage tasks like grasping in static environments via geometric and semantic attributes encoded in Gaussian properties. Multi-stage manipulation (Lu et al., 2024a; Shorinwa et al., 2024), where environmental dynamics demand real-time map updates, requires explicit modeling of dynamic adjustments (e.g., object motions and interactions), material compliance, etc.
@@ 5.2. Dynamic Scene Reconstruction
Dynamic scene reconstruction refers to the process of capturing and representing the three-dimensional structure and appearance of a scene that changes over time (Pumarola et al., 2021; Park et al., 2021b, a; Guo et al., 2023). This involves creating a digital model that accurately reflects the geometry, motion, and visual aspects of the objects in the scene as they evolve. Dynamic scene reconstruction is crucial in various applications, e.g., VR/AR, 3D animation, and autonomous driving (Zhou et al., 2024d; Yan et al., 2024b; Zhou et al., 2024a).
@@
The key to adapt 3D GS to dynamic scenes is the modeling of temporal dimension which allows for the representation of scenes that change over time. 3D GS based methods (Yang et al., 2024d, b; Lin et al., 2024b; Kratimenos et al., 2023; Das et al., 2024; Shaw et al., 2023; Wu et al., 2024b; Luiten et al., 2024; Huang et al., 2024d; Shao et al., 2024a; Yu et al., 2024a; Liang et al., 2023; Katsumata et al., 2023; Li et al., 2024i; Guo et al., 2024; Bae et al., 2024; Lei et al., 2024b; Wang et al., 2024c; Duan et al., 2024) for dynamic scene reconstruction can generally be divided into two main categories as discussed in Sec. 4.5 and Sec. 4.6. The first category utilizes additional fields like spatial MLPs or grids to model deformation (Sec. 4.6). For example, Yang et al. (Yang et al., 2024d) first proposed deformable 3D Gaussians tailored for dynamic scenes. These 3D Gaussians are learned in a canonical space and can be used to model spatial-temporal deformation with an implicit deformation field (implemented as an MLP). GaGS (Lu et al., 2024c) devised the voxelization of a set of Gaussian distributions, followed by the use of sparse convolutions to extract geometry-aware features, which are then utilized for deformation learning. On the other hand, the second category is based on the idea that scene changes can be encoded into the 3D Gaussian representation with a specially designed rendering process (Sec. 4.5). For instance, Luiten et al. (Luiten et al., 2024) introduced dynamic 3D Gaussians to model dynamic scenes by keeping the properties of 3D Gaussians unchanged over time while allowing their positions and orientations to change. Yang et al. (Yang et al., 2024b) designed a 4D Gaussian representation, where additional properties are used to represent 4D rotations and spherindrical harmonics, to approximate the spatial-temporal volume of scenes.
@@
While 3D GS advances dynamic scene reconstruction by modeling per-Gaussian deformations, its reliance on fine-grained primitives limits scalability and robustness. Current methods struggle to balance computational efficiency and precision: small-scale reconstructions unify dynamic and static elements but become intractable in large environments, often requiring manual priors to segment regions – a barrier in unstructured settings. Furthermore, the absence of object-level motion reasoning leads to artifacts and poor generalization over long sequences. Future work might prioritize object-centric frameworks that group Gaussians into persistent entities so as to model inherent motion disentanglement (dynamic vs. static).
@@ 5.3. Generation and Editing
Content generation and editing are two closely related capabilities in 3D creation pipelines. Generation aims to synthesize new assets or scenes from high-level conditions such as text and images, while editing focuses on modifying existing content in a controlled way. In this subsection, we concentrate on works that take 3D GS as the target representation. Please see the excellent survey (He et al., 2025) for more insights on the GS based generation and editing methods.
$\bullet$ Generation. Recent advances in GS based generation (Chen et al., 2024e; Tang et al., 2024b; Yi et al., 2024; Liang et al., 2024a; Liu et al., 2024c; Yang et al., 2023c; Zou et al., 2024; Ling et al., 2024; Ren et al., 2023; Yin et al., 2023; Zhang et al., 2023; Pan et al., 2024b; Xu et al., 2024a; Tang et al., 2024a; Yang et al., 2024a; Barthel et al., 2024; Jiang and Wang, 2024; Zhou et al., 2024c; Zhuo et al., 2024; Wu et al., 2024g; He et al., 2024; Yang and Wang, 2024; Kim et al., 2024; Feng et al., 2024a; Li et al., 2024b; Melas-Kyriazi et al., 2024; Li et al., 2024k; Zhang et al., 2024b; Mu et al., 2024; Lee et al., 2024c) can be roughly grouped along the underlying learning paradigm. Optimization-based methods follow the score-distillation paradigm: a powerful 2D or multi-view diffusion model is queried from many camera poses, and its gradients are back-propagated through the differentiable GS renderer to update Gaussian parameters. Early text- and image-to-3D pipelines such as DreamGaussian (Tang et al., 2024b) and GaussianDreamer (Yi et al., 2024) treat each scene independently and directly optimize positions, scales, colors, and opacities to match diffusion-model guidance while respecting RGB reconstruction losses. This per-scene fitting often yields high-fidelity geometry and appearance for object- and scene-level content, and can be naturally extended to dynamic or 4D generation with temporal regularization. However, it incurs heavy computational cost (minutes to hours per prompt) and is prone to Janus ambiguity, over-smoothing, or texture oversaturation unless additional regularizers, multi-view diffusion priors, or geometry-aware constraints are introduced. Reconstruction-based methods decompose the problem into two classical stages: multi-view generation followed by 3D reconstruction. Given a text or image prompt, a pre-trained multi-view or video diffusion model first synthesizes a small set of posed views (and sometimes depth or camera trajectories). A GS reconstructor such as LGM (Tang et al., 2024a) or VividDream (Lee et al., 2024c) then treats these images as supervision and solves a standard multi-view reconstruction problem to obtain Gaussians. This design reuses mature reconstruction pipelines and makes system-level engineering straightforward, but inherits the fundamental limitations of the multi-view generator: imperfect cross-view consistency can lead to noisy surface geometry, duplicated structures, and degraded textures, especially in occluded or textureless regions. Direct 3D generation methods aim to learn a feed-forward or diffusion model that outputs Gaussian fields directly from text or images, without an explicit intermediate image set (Mu et al., 2024; He et al., 2024; Zhang et al., 2024b; Pan et al., 2024a; Cai et al., 2025). Trained on large collections of synthetic or reconstructed data, these models amortize the reconstruction cost and naturally promote multi-view coherence, enabling near-real-time sampling at test time. The main challenges lie in the high-dimensional and highly structured output space (positions, anisotropic covariance, colors, opacity), the need for large, curated 3D training corpora, and maintaining controllability (e.g., disentangling geometry from appearance) while scaling to diverse objects and scenes. Scene-level versions further have to reason about layout, occlusions, and long-range dependencies. Beyond applications, these methods reflect a paradigm shift from per-scene optimization of Gaussians to amortized, feed-forward prediction (often with generative priors), enabling much lower test-time latency and better scalability (Sec. 7).
$\bullet$ Editing. Current editing works (Chen et al., 2024c; Fang et al., 2024; Huang and Yu, 2023; Ye et al., 2024; Cen et al., 2023; Zhou et al., 2024b; Lan et al., 2023; Huang et al., 2024d; Shao et al., 2024a; Yu et al., 2024a; Zhuang et al., 2024; Dou et al., 2024; Hu et al., 2024c; Palandra et al., 2024; Gu et al., 2024; Lyu et al., 2024; Liu et al., 2024f; Qiu et al., 2024; Zhang et al., 2024c; Wang et al., 2024e; Zhang et al., 2024g; Wu et al., 2024c) can be broadly divided into two families. The first leverages 2D image-editing models (e.g., text- or image-conditioned diffusion editors) to drive updates of 3D Gaussians. Representative systems such as GaussianEditor (Chen et al., 2024c; Fang et al., 2024; Palandra et al., 2024) repeatedly render GS views, apply 2D edits, and back-propagate the differences to Gaussian parameters, enabling text- or mask-driven appearance and geometry changes. Follow-up works improve spatial localization and multi-view consistency via attention, view correspondence, or iterative refinement (Wang et al., 2024e; Zhang et al., 2024g; Wu et al., 2024c), and specialize to style transfer or personalization (Zhuang et al., 2024; Zhang et al., 2024c). While intuitive and compatible with powerful 2D priors, these pipelines remain computationally heavy and can still suffer from unintended deformations due to weak 3D geometry priors. The second family exploits the explicit, point-based nature of 3D GS for direct manipulation. Semantic or feature-aware approaches enrich Gaussians with labels or embeddings so that users can select and edit subsets of points with text queries, masks, or keypoints (Cen et al., 2023; Ye et al., 2024; Hu et al., 2024c; Qiu et al., 2024; Huang et al., 2024d). Mesh-coupled methods provide stronger structural control. GaMeS (Waczyńska et al., 2024) introduces a hybrid “Gaussian Mesh Splatting” representation where each Gaussian is parameterized by the vertices of a mesh face; given either an input mesh or an estimated pseudo-mesh, classical mesh operations or animations induce coherent changes in Gaussian position, scale, and orientation. GaussianMesh (Gao et al., 2024) similarly defines Gaussians over an explicit mesh and ties the two via joint refinement: GS rendering guides adaptive face and Gaussian splitting, while mesh constraints regularize Gaussian shapes and support large, interactive deformations driven by mesh editing. However, this family faces several practical challenges: it relies on accurate semantic labels or mesh proxies, so errors in segmentation or remeshing can easily produce artifacts, and editing attributes (e.g., texture, geometry, and illumination) still requires careful regularization and cross-view alignment to preserve plausibility.